# Neural Bridge Processes

### v1

@@ Abstract
Learning stochastic functions from partially observed context-target pairs is a fundamental problem in probabilistic modeling. Traditional models like Gaussian Processes (GPs) face scalability issues with large datasets and assume Gaussianity, limiting their applicability. While Neural Processes (NPs) offer more flexibility, they struggle with capturing complex, multi-modal target distributions. Neural Diffusion Processes (NDPs) enhance expressivity through a learned diffusion process but rely solely on conditional signals in the denoising network, resulting in weak input coupling from an unconditional forward process and semantic mismatch at the diffusion endpoint. In this work, we propose Neural Bridge Processes (NBPs), a novel method for modeling stochastic functions where inputs $x$ act as dynamic anchors for the entire diffusion trajectory. By reformulating the forward kernel to explicitly depend on $x$, NBP enforces a constrained path that strictly terminates at the supervised target. This approach not only provides stronger gradient signals but also guarantees endpoint coherence. We validate NBPs on synthetic data, EEG signal regression and image regression tasks, achieving substantial improvements over baselines. These results underscore the effectiveness of DDPM-style bridge sampling in enhancing both performance and theoretical consistency for structured prediction tasks.
@@ 1 Introduction
Learning stochastic functions from partially observed context-target pairs is a fundamental problem in probabilistic modeling Rasmussen (2003); Garnelo et al. (2018b; a); Dutordoir et al. (2023); Franzese et al. (2023); Bonito et al. (2023); Mathieu et al. (2023); Chang et al. (2024); Dou et al. (2025); Hamad & Rosenbaum , playing a pivotal role in meta-learning Garnelo et al. (2018b; a), few-shot regression Kim et al. (2019), Bayesian optimization Dutordoir et al. (2023); Krishnamoorthy et al. (2023), and uncertainty-aware prediction tasks Chang et al. (2024); Dou et al. (2025); Hamad & Rosenbaum ; Requeima et al. (2024). Such problems require models that not only generalize well across different tasks but also provide calibrated uncertainty estimates, particularly under scarce or incomplete data conditions. Gaussian Processes (GPs) Rasmussen (2003) have traditionally dominated this area due to their analytical tractability and clear uncertainty quantification. However, GPs inherently assume Gaussianity and exhibit cubic computational complexity with respect to data size, severely limiting their applicability in scenarios involving large datasets or inherently non-Gaussian functional distributions Snelson & Ghahramani (2005); Titsias (2009); Xu & Zeng (2024); Xu et al. (2024).
@@
Neural Processes (NPs) Garnelo et al. (2018b; a); Kim et al. (2019); Louizos et al. (2019) have emerged as a compelling alternative, merging the flexibility of neural network models with the principled uncertainty quantification of stochastic processes. By parameterizing stochastic functions through neural architectures, NPs facilitate efficient inference and scalable learning, successfully tackling meta-learning and few-shot prediction tasks. However, standard NPs often struggle with limited expressivity and fail to capture complex multi-modal target distributions, motivating exploration into more powerful generative mechanisms.
NDPs Dutordoir et al. (2023) address this by modeling the input-output mapping as a learned diffusion process Ho et al. (2020), offering enhanced expressivity and sample diversity. Despite their promise in stochastic function modeling, NDPs rely on an unconditional forward process, which fundamentally limits the effectiveness of input supervision. Specifically, traditional NDPs treat inputs merely as conditional signals within the denoising network, passively injecting inputs during denoising without leveraging the temporal structure of diffusion. This results in weak input coupling and a semantic mismatch at the diffusion endpoint.
In this work, we introduce Neural Bridge Processes (NBP), a novel diffusion-based generative framework that explicitly integrates input supervision throughout the entire diffusion trajectory. Unlike traditional NDPs, which inject conditioning inputs passively during denoising, NBPs reformulate the forward diffusion kernel to dynamically anchor the process with inputs $x$, ensuring that the generated outputs remain coherently guided toward the desired targets. This is achieved through a principled bridge coefficient $\gamma_{t}$, which progressively strengthens the influence of $x$ as diffusion proceeds, enabling both strong gradient signals during training and guaranteed endpoint coherence. Additionally, NBPs incorporate a bridge correction term in the reverse process to maintain theoretical consistency between forward and reverse dynamics. Our approach provides a more structured and controllable generative path, leading to improved conditional generation accuracy and more faithful reconstructions, particularly in settings requiring strict adherence to input-output relationships. This idea shares conceptual similarities with recent advances in diffusion bridge modeling Zhou et al. (2023); Yue et al. (2023); Zheng et al. (2024); Li et al. (2023); Peluchetti (2023); He et al. (2024); Shi et al. (2023); Naderiparizi et al. (2025). However, our approach avoids the more complex and computationally intensive SDE-style diffusion bridges Song et al. (2020); Zhou et al. (2023), and instead extends the bridge concept to the DDPM  Ho et al. (2020) framework through SNR-aware functional modeling, making it significantly easier to deploy and integrate into existing architectures.
We validate NBPs on synthetic data and real time series data consisting of electroencephalogram (EGG) measurements Zhang et al. (1995) and image regression tasks, achieving substantial improvements over baseline NDPs. These results underscore the effectiveness of DDPM-style bridge sampling in enhancing both performance and theoretical consistency for structured prediction tasks.
@@
  • We introduce Neural Bridge Processes (NBPs), a new class of models for stochastic functions that introduces input-anchored diffusion trajectories via a principled bridge coefficient. This design ensures strong input supervision throughout the entire diffusion process, overcoming the weak coupling limitations of traditional NDPs and guaranteeing endpoint coherence.
  • We extend the bridge concept to the DDPM framework using an SNR- and path-aware formulation, thereby avoiding the deployment complexity associated with SDE-based diffusion bridges. This makes NBPs both theoretically consistent and practically efficient, enabling seamless integration into existing architectures.
  • We demonstrate the effectiveness of NBPs on synthetic data, EEG signal regression and image-based function regression benchmarks, achieving significant improvements in predictive accuracy and uncertainty calibration compared to state-of-the-art NDP baselines.
2 Background: Neural Processes
Neural Processes (NPs) Garnelo et al. (2018b; a); Kim et al. (2019); Louizos et al. (2019) combine the expressiveness of neural networks with the probabilistic reasoning of Gaussian Processes (GPs) Rasmussen (2003). While GPs offer principled uncertainty quantification, they suffer from poor scalability Snelson & Ghahramani (2005); Titsias (2009) and limited kernel flexibility Wilson et al. (2016); Liu et al. (2021). In contrast, Neural Networks (NNs) Schmidhuber (2015); Nielsen (2015) are highly flexible and scalable but lack inherent mechanisms for uncertainty modeling Blundell et al. (2015); Pearce et al. (2020); Gawlikowski et al. (2023). NPs address these limitations by modeling distributions over functions using a neural network-based framework. They approximate a stochastic process $F:X\to Y$ through finite-dimensional marginals, parameterized by a latent variable $z$ to capture global uncertainty. Given context observations $(x_{\mathbb{C}},y_{\mathbb{C}})$ and target inputs $x_{\mathbb{T}}$, NPs generate predictive distributions over $y_{\mathbb{T}}$ via a conditional latent model.
$p(y_{\mathbb{T}},z|x_{\mathbb{T}},x_{\mathbb{C}},y_{\mathbb{C}})=p(z|x_{\mathbb{C}},y_{\mathbb{C}})\prod_{i=1}^{|\mathbb{T}|}p(y_{\mathbb{T},i}|x_{\mathbb{T},i},z)$ (1)
$\log p(y_{\mathbb{T}}|x_{\mathbb{C}},y_{\mathbb{C}},x_{\mathbb{T}})\geq\mathbb{E}_{q(z|x_{\mathbb{C}},y_{\mathbb{C}})}\left[\sum_{i\in\mathbb{T}}\log p(y_{\mathbb{T},i}|z,x_{\mathbb{T},i})+\log\frac{p(z|x_{\mathbb{C}},y_{\mathbb{C}})}{q(z|x_{\mathbb{C}},y_{\mathbb{C}})}\right]$ (2)
where $q(z|x_{\mathbb{C}},y_{\mathbb{C}})$ is the variational posterior distribution parameterized by a neural network, and $p(z|x_{\mathbb{C}},y_{\mathbb{C}})$ is the conditional prior. Additional information can be seen in Appendix.
@@ 3.1 Problem Setup
We consider the standard meta-learning setting where a model observes a set of context points $x_{\mathbb{C}}=\{(x_{i},y_{i})\}_{i=1}^{N_{c}}$ and aims to predict the corresponding outputs $y_{\mathbb{T}}=\{y_{j}\}_{j=1}^{N_{t}}$ for a set of target inputs $x_{\mathbb{T}}=\{x_{j}\}_{j=1}^{N_{t}}$. Here, each task is assumed to be sampled from a distribution over functions, and the goal is to model the conditional distribution $p(y_{\mathbb{T}}|x_{\mathbb{C}},x_{\mathbb{T}})$.
3.2 Review: Neural Diffusion Processes (NDPs)
While Neural Processes (NPs) Garnelo et al. (2018b) effectively combine neural networks with stochastic processes for few-shot learning, their reliance on simple latent variable models limits their ability to capture complex, multimodal distributions. Neural Diffusion Processes (NDPs) Dutordoir et al. (2023) address this by introducing stochastic trajectories, modeling the mapping from inputs to outputs as a learned diffusion process Ho et al. (2020). This method significantly improves expressivity and sample diversity by leveraging the generative power of diffusion models Ho et al. (2020); Song et al. (2020); Dhariwal & Nichol (2021); Croitoru et al. (2023); Rombach et al. (2022); Podell et al. (2023); Peebles & Xie (2023), allowing for better modeling of complex distributions and more flexible conditional sampling.
Formally, given a function $f:\mathbb{R}^{D}\to\mathbb{R}$, an NDP learns a generative distribution over observed data pairs $(x,y)$, where inputs $x\in\mathbb{R}^{N\times D}$ and outputs $y=f(x)\in\mathbb{R}^{N}$. Unlike standard NPs, NDPs Dutordoir et al. (2023) do not explicitly require a partitioning into context and target sets during training; all points are jointly modeled. In supervised learning setting, the NDP modeling framework consists of two stochastic processes:
Forward Diffusion Process.
Reverse Process.
Training Objective.
NDPs employ a denoising score matching objective  Hyvärinen & Dayan (2005); Song et al. (2021); Huang et al. (2021), training the noise model $\epsilon_{\theta}$ by minimizing the discrepancy between predicted noise and actual noise $\epsilon\sim\mathcal{N}(0,I)$:
@@ 3.3 Motivation
Traditional Neural Diffusion Processes (NDPs) treat inputs merely as conditional signals fed into the denoising network (e.g., via cross-attention or concatenation). This implicit conditioning suffers from two critical drawbacks:
  • Weak coupling: The diffusion path is only loosely guided by inputs, as the forward process remains an unconditional Gaussian transition $q(y_{t}|y_{t-1})$.
  • Endpoint mismatch: The diffusion endpoint $y_{T}$ is arbitrary noise, bearing no semantic relationship to the input supervision $x$.
While NDPs have shown promise in generative modeling, their unconditional forward process fundamentally limits the efficacy of input supervision. Existing methods inject inputs passively during denoising, failing to exploit the temporal structure of diffusion. In this work, we propose Neural Bridge Processes (NBP), where inputs $x$ act as dynamic anchors for the entire diffusion trajectory. By reformulating the forward kernel to explicitly depend on $x$, NBP enforces a constrained path that strictly terminates at the supervised target. This approach not only provides stronger gradient signals but also guarantees endpoint coherence—a property unattainable by traditional NDPs.
@@ 3.4 Bridge Construction
We first consider the case where the input and output share the same dimensionality. Given a function $f:\mathbb{R}^{D}\to\mathbb{R}^{D}$, our Neural Bridge Processes (NBPs) model a generative distribution over observed data pairs $(x,y)$, with inputs $x\in\mathbb{R}^{N\times D}$ and outputs $y=f(x)\in\mathbb{R}^{N\times D}$. NBPs construct a diffusion bridge between arbitrary initial outputs $y_{0}=y$ and conditionally anchored endpoints characterized by the relationship between $\mathbb{E}[y_{T}]$ and $x$ via modified transition kernels.
To achieve this, we introduce a time-dependent coefficient $\gamma_{t}$ that explicitly controls the influence of the input $x$ on the forward diffusion process at each timestep $t$, thereby enabling explicit path supervision in contrast to standard DDPMs.
The forward transition kernel is defined as:
@@
The bridge coefficient $\gamma_{t}$ follows a principled design:
@@
Similar to DDPM, the forward process in this bridge-style diffusion model allows sampling $y_{t}$ at an arbitrary timestep $t$ in closed form:
@@
where the cumulative bridge coefficient is defined as
@@
The term $\gamma_{t}x$ in Equation (8) acts as a guiding force that progressively pulls the diffusion trajectory toward the target endpoint. This formulation ensures the following behavior:
  • Early Diffusion Phase ($t\ll T$): $\mathrm{SNR}_{t}\to+\infty\Rightarrow\gamma_{t}\to 0^{+}$. Process approximates standard diffusion:
    $q(y_{t}|y_{t-1},x)\approx\mathcal{N}(\sqrt{1-\beta_{t}}y_{t-1},\beta_{t}I)$ (12)
  • Bridge Convergence Phase ($t\to T$): $\mathrm{SNR}_{t}\to\mathrm{SNR}_{T}\Rightarrow\gamma_{t}\to 1$. The trajectory is increasingly guided toward the desired target to enforces endpoint attraction as shown in Equation (10):
    $\mathbb{E}[y_{T}|y_{0}]=\sqrt{\bar{\alpha}_{T}}y_{0}+\bar{\gamma}_{T}x\approx\bar{\gamma}_{T}x$ (13)
    where $\bar{\alpha}_{T}=\prod_{s=1}^{T}(1-\beta_{s})\approx 0$ , and $\bar{\gamma}_{T}=\left(\sum_{s=1}^{T}\gamma_{s}\sqrt{\frac{\bar{\alpha}_{T}}{\bar{\alpha}_{s}}}\right)$ is a constant.
This introduces:
1). Stronger gradient signals: The bridge term $\gamma_{t}x$ directly propagates input supervision to every timestep of the forward process by Equation (8) and (10), thereby introducing $x$ into the denoising network in a more theoretically consistent manner.
2). Coherent trajectory optimization: The entire path $y_{0:T}$ is trained to satisfy both data fidelity (to $y_{0}$) and endpoint matching (to $\mathbb{E}[y_{T}]=\bar{\gamma}_{T}x$).
Forward Process
Given the starting point $y_{0}$ and target endpoint $y_{T}$, the forward process progressively adds noise and enforces the bridge constraint through the transition kernel:
$q(y_{1:T}\mid y_{0},x)=\prod_{t=1}^{T}q(y_{t}\mid y_{t-1},x)$ (14)
The single-step transition kernel is defined in Equation (8). After $T$ diffusion steps, the distribution of the outputs converges to a Gaussian with mean $x$, i.e., $y_{T}\sim\mathcal{N}(\bar{\gamma}_{T}x,I)$.
@@
The reverse process employs a transition kernel that combines standard denoising with explicit bridge constraints to maintain consistency with the forward process:
@@
The mean function is reparameterized into two key components that balance denoising and bridge correction:
@@
where ${\alpha}_{t}=1-\beta_{t}$, and the bridge correction term $C_{t}(x)$ is derived as:
$C_{t}(x)=-\frac{\gamma_{t}}{\sqrt{1-\beta_{t}}}x.$ (17)
Proofs can be seen in Appendix. The role of the bridge correction term $C_{t}(x)$ is to ensure that the mean of the reverse process remains consistent with the bridge constraint imposed in the forward process. Specifically, in the forward process, the term $\gamma_{t}x$ injects information from the input $x$ into the diffusion trajectory. The reverse process compensates for this influence through $C_{t}(x)$.
Figure 1: The noise model architecture employed at each step of the Neural Bridge Processes. The greyed-out box highlights the bi-dimensional attention block.
@@
The training objective for NBPs minimizes the forward and reverse KL divergences, which is equivalent to minimizing the prediction error of the denoising network $\epsilon_{\theta}(y_{t},x,t)$ and the forward noise characterized by Equation (10).
$\mathcal{L}_{\theta}=\mathbb{E}_{t,x,y_{0},\epsilon}\left[\|\epsilon_{\theta}(y_{t},x,t)-\epsilon\|_{2}^{2}\right]\quad\text{with}\quad y_{t}=\sqrt{\bar{\alpha}_{t}}y_{0}+\bar{\gamma}_{t}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon.$ (18)
Proofs can be seen in Appendix. Here, $\epsilon\sim\mathcal{N}(0,I)$ represents the ground-truth noise added during the forward process. The reverse transition kernel’s dependence on both $y_{t}$ and $x$ ensures theoretical consistency, as the bridge variable $x$ is explicitly embedded in both the forward and reverse processes. This formulation guarantees that the learned reverse process remains properly coupled with the forward dynamics throughout the diffusion trajectory.

### v2

@@ Abstract
Learning stochastic functions from partially observed context-target pairs requires models that are expressive, uncertainty-aware, and strongly conditioned on inputs. Neural Diffusion Processes (NDPs) improve expressivity with denoising diffusion, but their forward process is input-independent; inputs only enter the reverse denoiser, so the noisy training states themselves do not encode the conditioning inputs. We propose Neural Bridge Processes (NBPs), which replace the unconditional forward kernel with an input-anchored bridge trajectory. When input and output dimensions differ, NBP learns an output-space anchor $a_{\psi}(x)=P_{\psi}(x)$, allowing coordinates or other inputs to guide the generative path without changing the denoising backbone. We show theoretically that process-level anchoring induces pathwise input distinguishability, injects information about $x$ into noisy states, and creates a direct gradient pathway unavailable to NDPs. Experiments on synthetic regression, EEG, CylinderFlow, and image regression show consistent improvements. Additional ablations show that the gains come from the full bridge construction with learned alignment, and that the same input-anchored path principle transfers to Flow Matching Neural Processes. These results suggest that bridge-anchored generative paths provide a general mechanism for strengthening conditional stochastic function modeling.
@@ 1 Introduction
Learning stochastic functions from partially observed context-target pairs is a fundamental problem in probabilistic modeling Rasmussen (2003); Garnelo et al. (2018b, a); Dutordoir et al. (2023); Franzese et al. (2023); Bonito et al. (2023); Mathieu et al. (2023); Chang et al. (2024); Dou et al. (2025); Hamad and Rosenbaum , with applications in meta-learning Garnelo et al. (2018b, a), few-shot regression Kim et al. (2019), Bayesian optimization Dutordoir et al. (2023); Krishnamoorthy et al. (2023), and uncertainty-aware prediction Chang et al. (2024); Dou et al. (2025); Hamad and Rosenbaum ; Requeima et al. (2024). These settings require models that generalize across tasks while remaining calibrated under scarce or incomplete observations. Gaussian Processes (GPs) Rasmussen (2003) have traditionally dominated this area due to their analytical tractability and clear uncertainty quantification. However, GPs inherently assume Gaussianity and exhibit cubic computational complexity with respect to data size, severely limiting their applicability in scenarios involving large datasets or inherently non-Gaussian functional distributions Snelson and Ghahramani (2005); Titsias (2009); Xu and Zeng (2024); Xu et al. (2024).
@@
Neural Processes (NPs) Garnelo et al. (2018b, a); Kim et al. (2019); Louizos et al. (2019) offer a scalable alternative by amortizing stochastic function prediction with neural architectures. However, standard NPs often struggle to capture complex multi-modal target distributions, motivating stronger generative mechanisms.
NDPs Dutordoir et al. (2023) address this by modeling the input-output mapping as a learned diffusion process Ho et al. (2020), offering enhanced expressivity and sample diversity. Despite their promise in stochastic function modeling, NDPs rely on an unconditional forward process. The input affects the reverse denoiser, but the noised state $y_{t}$ used during training is sampled from a distribution that does not depend on $x$. This creates a structural gap for conditional function modeling: input information is supplied only through the learned reverse network, rather than through the generative path itself.
In this work, we introduce Neural Bridge Processes (NBP), a diffusion-based generative framework that integrates input supervision throughout the diffusion trajectory. NBPs reformulate the forward diffusion kernel so that intermediate noisy states are dynamically anchored by the input through a bridge coefficient $\gamma_{t}$, together with a reverse correction term that keeps the reverse dynamics consistent with the modified forward process. When input and output dimensions differ, such as in coordinate-to-RGB image regression, we use a learned output-space anchor $a_{\psi}(x)=P_{\psi}(x)$ while leaving the denoising architecture unchanged; details and a visualization are provided in Appendix D.
NBP is related to conditional diffusion and diffusion bridge models, but differs in both the object being modeled and the location of conditioning. It learns distributions over functions and places the input anchor directly in the DDPM-style forward kernel, rather than only conditioning the denoiser or constructing a generic data-to-data bridge Zhou et al. (2023); Yue et al. (2023); Zheng et al. (2024); Li et al. (2023); Peluchetti (2023); He et al. (2024); Shi et al. (2023); Naderiparizi et al. (2025).
We provide a theoretical analysis showing that process-level anchoring gives NBP properties unavailable to standard NDPs: different inputs induce different forward marginals, the noisy state carries nonzero information about the input under a Gaussian input model, and the denoising objective receives an additional direct gradient pathway through $y_{t}$. We validate NBPs on synthetic data, electroencephalogram (EEG) regression Zhang et al. (1995), CylinderFlow, and image regression. Appendix H.2 isolates the source of the gains through component ablations, and Appendix H.3 shows that the same bridge idea transfers to Flow Matching Neural Processes (FlowNP) Hamad and Rosenbaum .
@@
  • We introduce Neural Bridge Processes (NBPs), a class of stochastic function models that use input-anchored diffusion trajectories to make the noisy training states explicitly depend on the conditioning inputs.
  • We formulate a DDPM-style bridge for function modeling, including a learned output-space anchor $a_{\psi}(x)=P_{\psi}(x)$ for dimension-mismatched inputs and outputs.
  • We provide theoretical analysis showing that process-level anchoring induces pathwise input distinguishability, input information in noisy states, and a direct gradient pathway for supervision.
  • We validate NBPs on synthetic regression, EEG, CylinderFlow, and image regression, and provide ablations showing that the gains come from the bridge construction rather than merely architecture size or denoiser capacity. We further show that the same bridge idea can improve a Flow Matching Neural Process, suggesting broader relevance for generative NP models.
2 Related Work
Neural Processes (NPs) Garnelo et al. (2018b, a); Kim et al. (2019); Louizos et al. (2019) learn distributions over functions from context-target observations, combining amortized neural prediction with uncertainty modeling. A broad line of work has improved NP expressivity through attention, convolutional structure, autoregressive modeling, graph structure, and other process-aware architectures Gordon et al. (2019); Foong et al. (2020); Singh et al. (2019); Nguyen and Grover (2022); Bruinsma et al. (2023); Hu et al. (2023). More recent methods incorporate stronger generative mechanisms into function modeling, including diffusion-based neural processes Dutordoir et al. (2023); Mathieu et al. (2023), score-based neural processes Dou et al. (2025), flow-based models Hamad and Rosenbaum , and related neural ODE formulations Chen et al. (2018); Norcliffe et al. (2021).
Our work is closest to Neural Diffusion Processes (NDPs) Dutordoir et al. (2023), which are already a conditional diffusion baseline for Neural Process-style function learning: they learn a DDPM-style reverse denoiser conditioned on the input and context observations. The key difference is that NDPs retain an unconditional forward diffusion process, so the noised training states are generated independently of the input. NBP keeps the same conditional denoising backbone but modifies the forward transition itself, so the input acts as a dynamic anchor throughout the diffusion trajectory. This process-level modification distinguishes NBP from standard conditional diffusion models Choi et al. (2021); Zhang et al. (2023); Zhu et al. (2023), where conditioning typically enters the reverse model while the forward noising process remains condition-independent.
NBP is also related to diffusion bridge models Zhou et al. (2023); Yue et al. (2023); Zheng et al. (2024); Li et al. (2023); Peluchetti (2023); He et al. (2024); Shi et al. (2023); Naderiparizi et al. (2025). Diffusion bridges have become popular in generative modeling, but existing DDBM-style methods are usually designed for data-space transport between paired endpoint distributions, such as image-to-image translation. This differs from the Neural Process setting, where the goal is to model stochastic functions from context-target observations and predict outputs at arbitrary target inputs. To the best of our knowledge, NBP is the first to introduce bridge-anchored generative paths into Neural Process-style function learning. The endpoint anchor is the function input, or a learned projection of it, and the bridge is used to strengthen conditional function prediction rather than generic domain translation. Appendix H.3 further compares this bridge principle with recent flow-based Neural Processes.
3 Preliminaries
Neural Processes (NPs) approximate a stochastic process $F:X\to Y$ through finite-dimensional marginals. Given context observations $(x_{\mathcal{C}},y_{\mathcal{C}})$ and target inputs $x_{\mathcal{T}}$, a latent-variable NP predicts target outputs by
$p(y_{\mathcal{T}},z|x_{\mathcal{T}},x_{\mathcal{C}},y_{\mathcal{C}})=p(z|x_{\mathcal{C}},y_{\mathcal{C}})\prod_{i=1}^{|\mathcal{T}|}p(y_{\mathcal{T},i}|x_{\mathcal{T},i},z)$ (1)
where $z$ captures uncertainty about the global structure of the underlying function. NDPs replace simple latent-variable decoders with a learned diffusion process over outputs, which improves expressivity but still uses an unconditional forward noising process.
3.1 Neural Diffusion Processes
Given clean outputs $y_{0}$, NDPs Dutordoir et al. (2023) define the standard DDPM forward process
$q(y_{1:T}\mid y_{0})=$ $\prod_{t=1}^{T}q(y_{t}\mid y_{t-1}),$ (2)
$q(y_{t}\mid y_{t-1})=$ $\mathcal{N}(y_{t};\sqrt{1-\beta_{t}}\,y_{t-1},\beta_{t}I).$
The terminal variable is therefore approximately standard Gaussian noise and does not encode the conditioning input. The reverse process is conditioned on $x$ through a denoising network:
where $p_{\theta}(y_{t-1}\mid y_{t},x)$ is parameterized by a noise predictor $\epsilon_{\theta}(y_{t},t,x)$. Training uses the usual denoising objective
$\mathcal{L}_{\theta}=$ $\mathbb{E}_{t,x,y_{0},\epsilon}\left[\left\|\epsilon-\epsilon_{\theta}(y_{t},t,x)\right\|_{2}^{2}\right],$ (4)
$\text{with}\quad y_{t}=$ $\sqrt{\bar{\alpha}_{t}}y_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon.$
This compact review highlights the key design choice that motivates NBP: in NDPs, $x$ affects the reverse denoiser but not the forward trajectory.
@@ 4.1 Problem Setup
We consider the standard meta-learning setting where a model observes a set of context points $x_{\mathcal{C}}=\{(x_{i},y_{i})\}_{i=1}^{N_{c}}$ and aims to predict the corresponding outputs $y_{\mathcal{T}}=\{y_{j}\}_{j=1}^{N_{t}}$ for a set of target inputs $x_{\mathcal{T}}=\{x_{j}\}_{j=1}^{N_{t}}$. Here, each task is assumed to be sampled from a distribution over functions, and the goal is to model the conditional distribution $p(y_{\mathcal{T}}|x_{\mathcal{C}},x_{\mathcal{T}},y_{\mathcal{C}})$. Throughout the method section, we use $x$ and $y$ to denote the full set of inputs and outputs for one task-level function, while $(x_{\mathcal{C}},y_{\mathcal{C}})$ and $(x_{\mathcal{T}},y_{\mathcal{T}})$ denote the context and target subsets used for conditional prediction. We use calligraphic symbols $\mathcal{C}$ and $\mathcal{T}$ for context and target index sets to avoid confusion with the complex numbers. The diffusion process is defined on the full task-level output set $y_{0:T}$ conditioned on the corresponding inputs $x$; the context–target partition is then instantiated during conditional sampling.
@@ 4.2 Motivation
Traditional Neural Diffusion Processes (NDPs) condition on inputs through the denoising network, for example by concatenation or attention. The forward process itself remains an unconditional Gaussian noising process, so the noisy state $y_{t}$ used for training is generated without direct dependence on $x$. By “indirect”, we mean that $x$ influences the model only through the learned denoiser $\epsilon_{\theta}(y_{t},x,t)$, while the distribution that produces $y_{t}$ is independent of $x$. Here, “temporal” refers to the ordered diffusion trajectory indexed by timestep $t$, not to an assumption that NP tasks themselves must have temporal dynamics. Once a diffusion-based NP model is adopted, there is an additional design choice: whether the input-output relation is imposed only through reverse-time conditioning, or also through the forward path used to generate training states. Standard NDPs choose the former.
NBP addresses this mismatch by making the input part of the forward transition. Instead of treating $x$ only as side information for the denoiser, we use $x$ as a dynamic anchor that changes the mean of every noising step. This coupling between $x$ and $y_{t}$ is therefore present throughout the trajectory rather than only in the denoiser. The resulting process remains DDPM-style and easy to train, but its trajectory is explicitly input-aligned in expectation.
To make this motivation concrete, we use Jacobian sensitivity as a diagnostic of effective input coupling:
$S(t)=\mathbb{E}\left[\left\|\frac{\partial\epsilon_{\theta}(y_{t},x,t)}{\partial x}\right\|_{F}\right].$ (5)
We compute this quantity on CelebA $32\times 32$ with context ratio 0.3 by averaging over 100 test images and multiple diffusion timesteps; implementation details are provided in Appendix H.4. Figure 1 shows that NDP has lower sensitivity to $x$ at early timesteps, whereas NBP has a larger measured sensitivity in the same region because $x$ also enters the forward state through the bridge term. We use this figure as supporting diagnostic evidence, not as a standalone proof of performance: the main quantitative evidence is provided by the likelihood, error, calibration, and ablation results.
Figure 1: Jacobian sensitivity across diffusion timesteps on CelebA $32\times 32$ with context ratio 0.3, averaged over 100 test images. NBP shows larger early-timestep sensitivity than NDP.
@@ 4.3 Bridge Construction
We first present the construction when the input and output share the same dimensionality. Given observed pairs $(x,y)$ with $x,y\in\mathbb{R}^{N\times D}$ and clean output $y_{0}=y$, NBP replaces the unconditional DDPM forward transition with
@@
The scalar $\gamma_{t}$ controls the strength of the input anchor. In this sense, the anchor is dynamic: the bridge term $\gamma_{t}x$ changes with the diffusion timestep and biases each intermediate state toward the input-conditioned endpoint with time-varying strength. We use the SNR-aware schedule
@@
When the input and output dimensions differ, we replace the raw input anchor $x$ by a learned output-space anchor $a_{\psi}(x)=P_{\psi}(x)\in\mathbb{R}^{N\times D_{y}}$ and apply the same bridge construction to $a_{\psi}(x)$. For example, in image regression, $x$ is a two-dimensional spatial coordinate while $y$ is an RGB value; the learned anchor network maps coordinates to the output channel space before they enter the bridge. Appendix D and Figure 3 provide the dimensional-alignment details. For readability, the equations below use $x$ to denote the aligned anchor, with the general case obtained by substituting $x$ with $a_{\psi}(x)$. This schedule keeps the early trajectory close to standard diffusion, where $\mathrm{SNR}_{t}$ is high and $\gamma_{t}$ is small, while increasing input attraction near the terminal region. Since $\gamma_{T}=\mathrm{SNR}_{T}/\mathrm{SNR}_{T}=1$, the last single-step transition has full input anchoring. This does not imply that the cumulative coefficient $\bar{\gamma}_{T}$ equals one; its scale is determined by the entire noise schedule. The following proposition gives the closed-form marginal used for training.
Proposition 4.1 (Input-anchored forward marginal).
For the transition in Eq. (6), the marginal distribution at any timestep $t$ is
@@
where
@@
The proof is provided in Appendix B.1. Proposition 4.1 shows that the terminal distribution is no longer input-independent. When $\bar{\alpha}_{T}\approx 0$, the terminal mean satisfies
$\mathbb{E}[y_{T}|y_{0},x]=\sqrt{\bar{\alpha}_{T}}y_{0}+\bar{\gamma}_{T}x\approx\bar{\gamma}_{T}x,$ (10)
while the terminal covariance is exactly $(1-\bar{\alpha}_{T})I$, which approaches $I$ only under a sufficiently long or strong noise schedule. Thus, the terminal state remains stochastic and schedule-dependent, but its mean becomes aligned with the input anchor in expectation.
@@
The reverse model keeps the DDPM denoising parameterization but adds the correction term induced by the input-dependent forward kernel:
@@
where $\alpha_{t}=1-\beta_{t}$, and
@@
Proposition 4.2 (Bridge posterior correction).
The posterior mean of $q(y_{t-1}\mid y_{t},y_{0},x)$ can be written in DDPM noise-prediction form plus the bridge correction
$C_{t}(x)=-\frac{\gamma_{t}}{\sqrt{1-\beta_{t}}}x=-\frac{\gamma_{t}}{\sqrt{\alpha_{t}}}x.$ (13)
The derivation is given in Appendix B.2. This correction is not an additional heuristic; it follows from the Gaussian posterior induced by the bridge transition in Eq. (6). Intuitively, the forward step injects $\gamma_{t}x$ into the mean, and the reverse mean must compensate for this deterministic input shift while denoising the stochastic component. This keeps the learned reverse process consistent with the bridge-anchored forward dynamics.
@@
Proposition 4.3 (Denoising objective).
Optimizing the variational bound for the bridge process reduces, up to the standard DDPM weighting, to the noise-prediction objective
$\mathcal{L}_{\theta}=$ $\mathbb{E}_{t,x,y_{0},\epsilon}\left[\|\epsilon_{\theta}(y_{t},x,t)-\epsilon\|_{2}^{2}\right]$ (14)
$\text{with}\quad y_{t}=$ $\sqrt{\bar{\alpha}_{t}}y_{0}+\bar{\gamma}_{t}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon.$
The proof is provided in Appendix B.3. Importantly, NBP does not require a new denoising architecture: the same noise predictor is trained with bridge-anchored samples from Proposition 4.1.